OpenStudy (anonymous):

Please help:) write the first five terms of the geometric sequences.. I suck at this okay so its... A1= 1, R= 1/3

OpenStudy (anonymous):

oh thanks i guess

OpenStudy (anonymous):

Just use the equation they show there an=(a1)r^(n-1) so a1=1*(1/3)^(1-1) then for a2 you would do: a2=a1r^(2-1) ect.

OpenStudy (anonymous):

okay thanks